# Energy Conditions and Quantum Information

### v1

@@ 4.5 Conformally invariant ANEC
A canonical example of spatially compact spacetimes is the static Einstein universe. As explained in Sec. 4.1, the conformal boundary $\partial M$ of $(d+1)$-dimensional AdS spacetime is globally the static Einstein universe in $d$-dimension, $\partial M\simeq{\mathbb{R}}\times S^{d-1}$. Let us consider a pair of two boundary points $p,q\in\partial M$ such that $p$ and $q\in J^{+}(\partial M)$ can be connected by an achronal null geodesic $\gamma$ with the tangent vector $k^{\mu}={\rm d}x^{\mu}/{\rm d}\lambda$ lying entirely in $\partial M$, and $q$ corresponds to the antipodal point of $p$ in spatial section $S^{d-1}$, as depicted in Figure 19. According to Gao-Wald theorem, under the bulk NEC, there is no bulk-shortcut connecting $p,q$ and $\gamma$ must be the fastest causal curve connecting the two point. Let $\lambda_{-}$ and $\lambda_{+}$ be the affine parameter values of $\gamma$ at the point $p$ and $q$, respectively, and $\eta$ be the magnitude of the boundary Jacobi field along $\gamma$, which vanishes at $p,q$, i.e., $\eta(\lambda_{\pm})=0$ but is non-vanishing at any other points between $p$ and $q$, due to the achronality of $\gamma$ (Proposition 2.5.1). On this curved boundary, we can evaluate the expectation value of the energy-momentum tensor $\langle T_{\mu\nu}\rangle$ by using the holographic formula (97). Note that when the boundary dimension $d$ is odd, $X_{\mu\nu}$ in (97) identically vanishes, while when $d$ is even, $X_{\mu\nu}$ must be taken into account. Then, applying Gao-Wald theorem, we find the following inequality: {itembox}[l]Definition 4.5.1 (Holographic weighted ANEC)
$\int_{\lambda_{-}}^{\lambda_{+}}\!\!\ {\rm d}\lambda\>\eta^{d}\langle T_{\mu\nu}\rangle k^{\mu}k^{\nu}\geqslant C\,.$ (104)
@@ 5 Summary and discussions
As we have seen throughout this paper, locally defined energy conditions are often violated by quantum effects, such as vacuum fluctuations or the Casimir effect. It is especially noteworthy that the ANEC is not a universal feature of quantum field theories. Yet, in holographic theories with consistent bulk duals, the ANEC is restored as a consequence of deeper geometric and information-theoretic principles. Perhaps most strikingly, this result reverses the typical logic of holography: rather than using boundary data to constrain the bulk, bulk causality is used to derive nontrivial dynamical constraints on the boundary theory. The argument elegantly ties together extremal surfaces, entanglement entropy, and modular flow to derive a physical energy condition from first principles.

### v2

@@ 4.5 Conformally invariant ANEC
A canonical example of spatially compact spacetimes is the static Einstein universe. As explained in Sec. 4.1, the conformal boundary $\partial M$ of $(d+1)$-dimensional AdS spacetime is globally the static Einstein universe in $d$-dimension, $\partial M\simeq{\mathbb{R}}\times S^{d-1}$. Let us consider a pair of boundary points $p,q\in\partial M$. Fix $p\in\partial M$ and choose $q\in J^{+}(p,{\partial M})$ so that there exists an achronal boundary null geodesic $\gamma\subset\partial M$ with tangent $k^{\mu}=\mathrm{d}x^{\mu}/\mathrm{d}\lambda$ connecting $p$ to $q$; take $q$ to be the antipodal point of $p$ on $S^{d-1}$ (Fig. 19). According to Gao-Wald theorem, under the bulk NEC, there is no bulk-shortcut connecting $p,q$ and $\gamma$ must be the fastest causal curve connecting the two point. Let $\lambda_{-}$ and $\lambda_{+}$ be the affine parameter values of $\gamma$ at the point $p$ and $q$, respectively, and $\eta$ be the magnitude of the boundary Jacobi field along $\gamma$, which vanishes at $p,q$, i.e., $\eta(\lambda_{\pm})=0$ but is non-vanishing at any other points between $p$ and $q$, due to the achronality of $\gamma$ (Proposition 2.5.1). On this curved boundary, we can evaluate the expectation value of the energy-momentum tensor $\langle T_{\mu\nu}\rangle$ by using the holographic formula (99). Note that when the boundary dimension $d$ is odd, $X_{\mu\nu}$ in (99) identically vanishes, while when $d$ is even, $X_{\mu\nu}$ must be taken into account. Then, applying Gao-Wald theorem, we find the following inequality:
  Definition 4.5.1 (Holographic weighted ANEC)   $\int_{\lambda_{-}}^{\lambda_{+}}\!\!\ {\rm d}\lambda\>\eta^{d}\langle T_{\mu\nu}\rangle k^{\mu}k^{\nu}\geqslant C\,.$ (106)     where $\eta$ is the magnitude of the Jacobi field of the null geodesic congruence $\gamma$, representing the separation of points between the two adjacent null geodesics on $\partial M$.
@@ 5 Summary and discussions
As we have seen throughout this paper, locally defined energy conditions are often violated by quantum effects, such as vacuum fluctuations or the Casimir effect. It is especially noteworthy that the ANEC is not a universal feature of quantum field theories. Yet, in holographic theories with consistent bulk duals, the ANEC is restored as a consequence of deeper geometric and information-theoretic principles. Perhaps most strikingly, this result reverses the typical logic of holography: rather than using boundary data to constrain the bulk, bulk causality is used to derive nontrivial dynamical constraints on the boundary theory. The argument elegantly ties together extremal surfaces, entanglement entropy to derive a physical energy condition from first principles.